Use taylors series method to find approximate Value of y at x =0.2,given that dy/dx =2y+3ex,y(0)=0 comapre the result by solving the differential equation


1
Expert's answer
2022-02-08T15:25:29-0500

We are given; "\displaystyle
y\prime=2y+3e^x,\ y(0)=0,\ h=0.1" (since step length is not given, I assume a step length of "0.1"), "\displaystyle
x_0=0,\text{and } y_0=0".

Using Taylor's series method, we are to find "y(0.2)."

Now, differentiating the given DE successively yields;

"\displaystyle
y\prime=2y+3e^x\\
y\prime\prime=2y\prime+3e^x\\
y\prime\prime\prime=2y\prime\prime+3e^x\\"


Now, substituting yields;

"\displaystyle
y_0\prime=2y_0+3e^{x_0}=3\\
y_0\prime\prime=2y_0\prime+3e^{x_0}=9\\
y_0\prime\prime\prime=2y_0\prime\prime+3e^{x_0}=21\\"


Putting these values into the Taylor's series yields;

"\displaystyle
y(0.1)=y_1=y_0+hy_0\prime+\frac{h^2}{2!}y_0\prime\prime+\frac{h^3}{3!}y_0\prime\prime\prime+\cdots"

"\displaystyle
=0+0.1(3)+\frac{0.1^2}{2}\times9+\frac{0.1^3}{6}\times21+\cdots=0.3485"


Next, "x_1=0.1", and "y_1=0.3485". Thus substituting yields;

"\displaystyle
y_1\prime=2y_1+3e^{x_1}=4.012512754\\
y_1\prime\prime=2y_1\prime+3e^{x_1}=11.34053826\\
y_1\prime\prime\prime=2y_1\prime\prime+3e^{x_1}=25.99658928\\"

Putting these values into the Taylor's series yields;

"\displaystyle
y(0.2)=y_2=y_1+hy_1\prime+\frac{h^2}{2!}y_1\prime\prime+\frac{h^3}{3!}y_1\prime\prime\prime+\cdots"

"\displaystyle
=0.3485+0.1(4.012512754)+\frac{0.1^2}{2}\times11.34053826"

"\displaystyle
+\frac{0.1^3}{6}\times25.99658928+\cdots=0.8107867316"

Thus, the approximate value of "y(0.2)" using Taylor's series is;

"y(0.2)=0.8107867316"


Solving directly yields;

"\displaystyle
y\prime=2y+3e^x\Rightarrow y\prime-2y=3e^x"

Integrating Factor is;

"\displaystyle
e^{\int(-2)\ dx}=e^{-2x}"

"\displaystyle""\displaystyle
(y\prime-2y)e^{-2x}=3e^x\times e^{-2x}\\
\Rightarrow \frac{d}{dx}(ye^{-2x})=3e^{-x}\\
\Rightarrow ye^{-2x}=\int3e^{-x}\ dx=3\int e^{-x}\ dx=-3e^{-x}+c", where c is an arbitrary constant

"\displaystyle
\Rightarrow y=ce^{2x}-3e^{x}"

Since "\displaystyle
y(0)=0, \text{ then }c=3\Rightarrow y=3e^{2x}-3e^{x}\\"

Now,

"\displaystyle
y(0.2)=3e^{2\times0.2}-3e^{0.2}=3e^{0.4}-3e^{0.2}=0.8112658184"